OpenStudy (anonymous):

The difference between the circumference and the radius of a circle is 37 cm.The area of the circle is........................

OpenStudy (anonymous):

\[Circumference = 2 \pi .r\] \[2 \pi r - r = 37\] \[r(2 \pi - 1) = 37\] Solve for r here by putting pi = 3.14 Then the value of r you will put into the formula below to find the area: \[Area = \pi.r^2\]

OpenStudy (anonymous):

r(2*pi-1)=37 r(2*3.14-1)=37 r*5.28=37 is the process correct?

OpenStudy (anonymous):

Yes you are going right.. find r from it..

OpenStudy (anonymous):

Divide by 5.28 both the sides.. can you do it or not?

OpenStudy (anonymous):

i try.

OpenStudy (anonymous):

Take your time..

OpenStudy (anonymous):

7 ?

OpenStudy (anonymous):

radius=7?

OpenStudy (anonymous):

Yes you got it.. now for simplicity put in area the value of pi as 22/7 and not 3.14 Getting??

OpenStudy (anonymous):

yeah!

OpenStudy (anonymous):

154cmsq.

OpenStudy (anonymous):

Yes dear...

OpenStudy (anonymous):

hey if u don't have any prob. can i ask u one more ques.?

OpenStudy (anonymous):

Sure I will see what can I do in that..

OpenStudy (anonymous):

The inner and the outer radii of a cylindrical pipe are 5cm and 4cm respectively.Find the area of cross section of the pipe(Take pi=3.14).

OpenStudy (anonymous):

wat happen? is it hard?

OpenStudy (anonymous):

I don't know the formula for area of cross section of pipe.. Do you know?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Try this one if I am thinking right: \[Area = \pi .(R - r)^2\] R = Outer Radius.. r = Inner Radius..

OpenStudy (anonymous):

Able to do or not??

OpenStudy (anonymous):

pi*(5-4)sq. like this?

OpenStudy (anonymous):

Yes solve it and check if the answer is coming right or not?

OpenStudy (anonymous):

Area simply becomes equal to pi, I think...

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Is the answer right?

OpenStudy (anonymous):

i am asking u!

OpenStudy (anonymous):

Don't you have the answers with you??

OpenStudy (anonymous):

i am making an answer key!!!

OpenStudy (anonymous):

So, you should post this question so that the answer can be confirmed..

OpenStudy (anonymous):

i've already bumped it!